OpenStudy (anonymous):

SIMPLIFY (4r^9 OVER r^6s^4)^2

OpenStudy (anonymous):

\[(\frac{4r^9}{r^6s^4})^2\] \[=\frac{4^4r^{18}}{r^{12}s^8}\]

OpenStudy (anonymous):

\[=\frac{16r^6}{s^8}\]